The perimeter of a rectangle is 30 inches , and its area is 54 square inches. Find the length of the longest side of the rectangle
Answer by stanbon(75887)

The perimeter of a rectangle is 30 inches , and its area is 54 square inches. Find the length of the longest side of the rectangle.
------
P = 2(Length + Width)
------
Length + Width = 15 inches
Length * Width = 54 in^2
---------------------------------
L + W = 15
L*W = 54
-----
Substitute for "L" and solve for "W":
(15-W)*W = 54
-----
-W^2 + 15W - 54 = 0
--------------------------
W^2 - 15W + 54 = 0
----
W = 6 or W = 9
-----
If Width = 6, Length = 9
If Width = 9, Length = 6
-------------
Cheers,
Stan H.
--------------